Let X be a Bernoulli random variable (Hint: Special case of Binomial

Distribution)

a. Compute E(X2)

b. Show that V(X) is p(1 − p).

c. Compute E(X79)

1
Expert's answer
2021-09-28T08:39:47-0400

If X is A Bernoulli random variable, then its p. d. f is given by

"P(X=x)=p^{x}(1-p)^{1-x}, x=0, 1"

First, determine P(X=x) when x=0 and x=1.

So,

P(X=0)=p0(1-p)1-0=1*(1-p)=1-p

P(X=1)=p1(1-p)1-1=p*1=p

a.

E(X2)="\displaystyle\sum_{x=0}^{1}(\def\foo{x^2}\foo)*P(X=x)"


=02*P(X=0)+12*P(X=1)

=0*(1-p)+1*(p)=p

Therefore, E(x2)=p

b. Variance of random variable X is defined by,

V(X)=E(X2)-(E(X))2

We need to determine E(X) which is given as,

E(X)="\displaystyle\sum_{x=0}^{1}(x)*P(X=x)"

=0*P(X=0)+1*P(X=1)

=0*(1-p)+1*(p)=p

Therefore, V(X)=p-p2

Factoring gives p(1-p)

Hence V(X)=p(1-p) as given.

c.

E(X79) ="\displaystyle\sum_{x=0}^{1}(x^{79})*P(X=x)"


=(079)*P(X=0)+(179)*P(X=1)

=0*(1-p)+1*p

=p